# Task-Unaware Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

### v1

@@ Abstract
Real-world environments require robots to continuously acquire new skills while retaining previously learned abilities, all without the need for clearly defined task boundaries. Storing all past data to prevent forgetting is impractical due to storage and privacy concerns. To address this, we propose a method that efficiently restores a robot’s proficiency in previously learned tasks over its lifespan. Using an Episodic Memory (EM), our approach enables experience replay during training and retrieval during testing for local fine-tuning, allowing rapid adaptation to previously encountered problems without explicit task identifiers. Additionally, we introduce a selective weighting mechanism that emphasizes the most challenging segments of retrieved demonstrations, focusing local adaptation where it is most needed. This framework offers a scalable solution for lifelong learning in dynamic, task-unaware environments, combining retrieval-based adaptation with selective weighting to enhance robot performance in open-ended scenarios.
@@ 1 Introduction
Figure 1: Method Overview. Our approach addresses the challenge of lifelong learning without distinct task boundaries. To emulate human learning patterns, we propose a method consisting of three phases: learning, reviewing, and testing. In the learning phase, the robot is exposed to various demonstrations, storing a subset of this data as episodic memory $\mathcal{M}$. During the reviewing phase, the method retrieves the most relevant data to fine-tune the policy network, enhancing performance in the final testing phase.
Lifelong learning seeks to endow neural networks with the ability to continually acquire new skills while retaining previously learned knowledge. This balance between stability and plasticity is crucial as models face sequences of tasks over time. While significant progress has been made in applying lifelong learning to domains such as computer vision (Huang et al., 2024; Du et al., 2024) and natural language processing (Shi et al., 2024; Razdaibiedina et al., 2023), the challenges are more pronounced in robotics. Robots are expected to adaptively learn and solve unseen tasks throughout their operational lifespan (Thrun & Mitchell, 1995). Their interactions with dynamic environments introduce complexities absent in static data domains; a single misstep in task execution can result in complete failure. Moreover, robotics is constrained by limited data availability due to the expense and complexity of real-world interactions (Zhu et al., 2022; Du et al., 2023). These factors not only intensify the difficulty of continual learning in robotics but also demand more robust lifelong learning capabilities.
Existing methods for lifelong robot learning typically require robots to learn a sequence of tasks, each distinguished by domain, scenario, scene, or task goals (Liu et al., 2024; Yang et al., 2022; Wan et al., 2024; Parakh et al., 2024). In those settings, robots often depend on specific task identifications with clear boundaries — usually provided as task IDs or explicit descriptions — to specify the task they are working on (Liu et al., 2023). However, in dynamic real-world settings, it is impractical to predefine tasks or assign specific IDs, as robots are likely to encounter a vast array of unpredictable situations, with tasks that may be subdivided into smaller components of varying granularity. Therefore, approaches that rely on specific task identifications with clear boundaries are unrealistic and unscalable (Koh et al., 2021).
To address these challenges, we propose a novel task-unaware lifelong robot learning framework with visuomotor policies, utilizing vision perceptions as well as diverse paraphrased language descriptions. This framework enables robots to continually learn and adapt without explicit task IDs. We employ our method in manipulation scenarios based on the LIBERO benchmark (Liu et al., 2024). Our approach leverages pre-trained models to generate consistent embeddings across different tasks and training phases, thereby mitigating the embedding drift that often occurs in sequential learning scenarios (Liu et al., 2023; Kawaharazuka et al., 2024). We adopt Experience Replay (ER) baseline (de Masson D’Autume et al., 2019) to rehearse samples from previous tasks, helping to maintain learned skills and reduce forgetting.
Despite these measures, some degree of forgetting remains inevitable due to the multitasking nature of lifelong learning and the robot’s limited access to previous demonstrations. Drawing inspiration from human learning processes — where individuals revisit tasks they once knew but have forgotten details — we introduce an efficient local adaptation mechanism. Humans often perform quick reviews using limited resources and try to retrieve memory to rebuild their knowledge, allowing them to efficiently regain proficiency without relearning all aspects of the task (Sara, 2000). Similarly, our mechanism enables the robot to locally adapt to previously encountered problems rapidly and regain skills through fast fine-tuning, using the same episodic memory employed for experience replay during training.
Given the indistinct task boundaries, we leverage retrieval-based mechanisms (Du et al., 2023; van Dijk et al., 2024; de Masson D’Autume et al., 2019) to retrieve data most similar to the current task based on vision and language input similarities. To adapt the model effectively — especially focusing on the most challenging phases where the robot’s performance deviates — we first perform a few episodes of rollouts to obtain ”feedback” on the model’s performance before local adaptation: these rollouts are then used for automatic selective weighting by comparing them with the retrieved demonstrations without human intervention (Spencer et al., 2022; Mandlekar et al., 2020). The weighted samples facilitate the local adaptation phase, thereby improving performance.
In summary, the key contributions of our solution are:
  • Retrieval-Based Adaptation for Blurred Task Boundaries: During testing, relevant past demonstrations are retrieved from episodic memory to adapt the neural network locally, enabling the robot to quickly regain proficiency on previously encountered tasks without relying on explicit task boundaries.
@@
  • Selective Weighting Mechanism: A weighting mechanism emphasizes the most challenging segments of the retrieved demonstrations, optimizing real-time adaptation.
  • Paradigm for Memory-Based Lifelong Robot Learning: We demonstrate that our approach can be applied to different memory-based robotic lifelong learning algorithms during test time, serving as a paradigm for skill restoration.
This framework allows robots to continually learn and adapt in dynamic environments without requiring predefined task identifiers or boundaries, making it highly practical and scalable for real-world applications. By combining retrieval-based adaptation with selective weighting, our method offers a robust solution to the challenges of lifelong robot learning in open-ended settings.
2 Related Work
@@ 2.1 Lifelong Robot Learning
A key challenge in lifelong robot learning is catastrophic forgetting, where learning new tasks adversely affects performance on previously learned tasks (Parisi et al., 2019). Traditional lifelong learning methods often rely on explicit task identifiers or clear task boundaries to structure the learning process (Wan et al., 2024; Xie & Finn, 2022), such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017) and PackNet (Mallya & Lazebnik, 2018). In real-world robotic applications, robots operate in dynamic environments where tasks are not clearly segmented, making explicit task identifiers impractical (Kim et al., 2024).
Recent efforts in lifelong reinforcement learning (Xie & Finn, 2022) and areas like task and motion planning (Mendez-Mendez et al., 2023), online object model reconstruction (Lu et al., 2022), interactive instruction-following agents (Kim et al., 2024), multi-task learning (Wang et al., 2022), interactive imitation learning (Spahn et al., 2024), and SLAM (Yin et al., 2023; Gao et al., 2022; Vödisch et al., 2022) show progress. Robot manipulation skills also evolve through interactions, aiding adaption when task executions fail (Parakh et al., 2024). Memory-based algorithms and selective weighting (Sun et al., 2022; Koh et al., 2021; Shim et al., 2021; Aljundi et al., 2019) enhance learning by prioritizing informative samples. Replay buffer methods (He et al., 2020; Mai et al., 2021; Caccia et al., 2021) have demonstrated success. However, there is still a lack of progress in settings where the model is unaware of task boundaries during both training and inference (Lee et al., 2020; Chen et al., 2020; Ardywibowo et al., 2022).
A benchmark for lifelong robot learning, particularly focusing on manipulation tasks, has been introduced in LIBERO (Liu et al., 2024). Methods such as TAIL (Liu et al., 2023) rely on specific task identifiers, which can be limiting in dynamic environments, while Lotus (Wan et al., 2024) involves a pretraining phase to establish an initial skill set, providing a foundation for further continuous learning.
2.2 Task-unaware Continual Learning
Despite the success of continual learning with clearly labeled task sequences, there still remains a gap in progress within settings where the model is unaware of task boundaries both in training and inference, an online situation more reflective of real-world scenarios. Many attempts (Lee et al., 2020; Chen et al., 2020; Ardywibowo et al., 2022) focus on learning specialized parameters using expanding network structures. Memory-based algorithms remain effective by prioritizing informative samples (Sun et al., 2022), removing less important training samples (Koh et al., 2021), improving decision boundaries (Shim et al., 2021), and increasing gradient diversity (Aljundi et al., 2019). Methods aiming to exploit replay buffer (He et al., 2020; Mai et al., 2021; Caccia et al., 2021) have also demonstrated notable success.
2.3 Information Retrieval for Robotics
Information retrieval techniques have been used to optimize robotic behaviors by retrieving relevant actions from memory in novel tasks (Du et al., 2023). For example, path following based on image retrieval improves visual navigation (van Dijk et al., 2024), and incremental learning helps humanoid robots adapt to new environments by recalling past behaviors (Bärmann et al., 2023). Retrieval has also enabled skill transfer from videos (Papagiannis et al., 2024) and affordance transfer for zero-shot manipulation (Kuang et al., 2024), allowing robots to manipulate objects without prior training.
@@ 2.4 Robot Learning with Adaptation
Robots have learned to adapt to dynamic environments, such as agile flight in strong winds (O’Connell et al., 2022) and quadruped locomotion adaptation through test-time search (Peng et al., 2020). Meta-learning allows fast adaptation to new tasks (Kaushik et al., 2020; Nagabandi et al., 2018), while efficient adaptation techniques enable robots to generalize from limited data (Julian et al., 2020). Despite the success of retrieval-based adaptation and selective weighting, combining these methods for lifelong learning in open-ended environments remains an open challenge.
@@
Figure 2: Policy Backbone Architecture used in Training and Testing. We input various data modalities into the system, including demonstration images, language descriptions, and the robot arm’s proprioceptive input (joint and gripper states). Pretrained R3M (Nair et al., 2022) and SentenceSimilarity (2024) models process the image and language data respectively. Along with the proprioceptive states processed by an MLP, the embeddings are concatenated and passed through a Transformer to generate temporal embeddings. A GMM (Gaussian Mixture Model) is then used as the policy head to sample actions for the robot. Throughout both training and testing, we utilize episodic memory to store a subset of demonstrations gathered throughout the training process.
3 Preliminary
Our robot utilizes a visuomotor policy learned through behavior cloning to execute manipulation tasks by mapping sensory inputs and task descriptions to motor actions. In a task-unaware lifelong learning setting, we adopt a continual learning framework where task boundaries are blurred by employing multiple paraphrased descriptions to define task goals, rather than relying on explicit task identifiers. This approach enhances the policy’s ability to generalize across varied instructions and tasks.
@@ 3 Preliminary
The policy is trained by minimizing the discrepancy between the predicted actions and the expert actions derived from demonstrations. Specifically, we optimize the following loss function across a sequence of tasks $\{\mathcal{T}_{k}\}$ with corresponding demonstrations $\mathcal{D}_{k}=\{\tau^{k}_{1},\ldots,\tau^{k}_{N}\}$. Notably, $\mathcal{D}_{k}$ is not fully accessible for $k<K$ due to the use of experience replay data from Episodic Memory $\mathcal{M}$:
$\theta^{*}=\arg\min_{\theta}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(o_{t},a_{t})\sim\mathcal{D}_{k},\,g\sim G_{k}}\left[\sum_{t=0}^{l_{k}}\mathcal{L}\left(\pi_{\theta}(o_{\leq t},g),\,a_{t}\right)\right],$ (1)
@@
where $\theta$ denotes the model parameters, $l_{k}$ represents the number of samples for task $k$, $o_{\leq t}$ denotes the sequence of observations up to time $t$ in demonstration $n$ (i.e., $o_{\leq t}=(o_{0},o_{1},\ldots,o_{t})$), and $a_{t}$ is the expert action at time $t$. The set $G_{k}$ comprises paraphrased goal descriptions for task $\mathcal{T}_{k}$, with $g$ being a sampled goal description from $G_{k}$. The policy output, $\pi_{\theta}(o_{\leq t},g)$, is conditioned on both the observation sequence and the goal description.
@@
By optimizing this objective function, the policy effectively continues learning new tasks and skills in its life span, without the need for explicit task labels, thereby facilitating robust and adaptable task-unaware continual learning.
@@
4 Retrieval-based Weighted Local Adaptation for Lifelong Robot Learning
@@
In this section, we outline our proposed method depicted in Figure 1. To effectively interact with complex physical environments, the network integrates multiple input modalities, including visual inputs from workspace and wrist cameras, proprioceptive inputs of joint and gripper states, and paraphrased task descriptions.
@@
Instead of training all modules jointly in an end-to-end manner, we employ pretrained visual and language encoders that leverage prior semantic knowledge. Pretrained encoders enhance performance on downstream manipulation tasks (Liu et al., 2023) and are well-suited to differentiate between various scenarios and tasks without relying on explicit task identifiers or clear task boundaries. Their consistent representations when new tasks continue to come is essential for managing multitask problems and retrieving relevant data to support our proposed local adaptation during test time.
@@
When learning new tasks, the robot preserves previously acquired skills by replaying prior manipulation demonstrations stored in an episodic memory $\mathcal{M}$, which contains a small subset of previous task demonstrations (Chaudhry et al., 2019). Trained with the combined data from the latest scenarios and episodic memory $\mathcal{M}$, the model can acquire new skills while mitigating catastrophic forgetting of old tasks, thereby maintaining a balance between stability and plasticity (Wang et al., 2024). Figure 2 illustrates the network architecture, and implementation details are provided in Section A.2.

### v3

# Task-free Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

@@ Abstract
A fundamental objective in intelligent robotics is to move towards lifelong learning robot that can learn and adapt to unseen scenarios over time. However, continually learning new tasks would introduce catastrophic forgetting problems due to data distribution shifts. To mitigate this, we store a subset of data from previous tasks and utilize it in two manners: leveraging experience replay to retain learned skills and applying a novel Retrieval-based Local Adaptation technique to restore relevant knowledge. Since a lifelong learning robot must operate in task-free scenarios, where task IDs and even boundaries are not available, our method performs effectively without relying on such information. We also incorporate a selective weighting mechanism to focus on the most ”forgotten” skill segment, ensuring effective knowledge restoration. Experimental results across diverse manipulation tasks demonstrate that our framework provides a scalable paradigm for lifelong learning, enhancing robot performance in open-ended, task-free scenarios.
@@ 1 Introduction
Figure 1: Method Overview. Our approach addresses the challenge of lifelong learning without relying on task boundaries or IDs. To emulate human learning patterns, we propose a method consisting of three phases: Lifelong Learning, Reviewing, and Testing. In the Lifelong Learning phase, the robot is exposed to various demonstrations, storing a subset of the data in a storage memory $\mathcal{M}$. During the Reviewing phase before policy deployment, the method retrieves the most relevant data to locally adapt the policy network, enhancing performance in the deployment scenario.
Significant progress has been made in applying lifelong learning to domains such as computer vision (Huang et al., 2024; Du et al., 2024) and natural language processing (Shi et al., 2024; Razdaibiedina et al., 2023). However, extending lifelong learning to robotics poses additional challenges, as robots must interact with the real world under sequential decision-making constraints. Besides, the high cost and complexity of physical interactions (Zhu et al., 2022; Du et al., 2023) limit data availability, emphasizing the need for more effective strategies to mitigate catastrophic forgetting and maintain performance over robots’ lifelong run (Thrun & Mitchell, 1995). Furthermore, in realistic and scalable scenarios, lifelong learning robots must operate in a task-free setting, where they are not provided with specific task IDs or boundaries for each new task. This further complicates the challenge, as robots must continually adapt to new tasks without prior knowledge of task distinctions.
In practical lifelong robot learning settings (Liu et al., 2024, 2023), a robot learns a series of tasks in a sequential manner. Lacking future data or environments, while also having only partial or no access to past data due to storage or security restrictions, leads to significant data distribution shifts and hinders stable lifelong learning. Although numerous approaches address this catastrophic forgetting problem (Wang et al., 2024), many rely on task boundaries or explicit task identifiers to consolidate knowledge, limiting their scalability in open-ended real-world scenarios (Koh et al., 2021).
To tackle these issues, we propose a task-free memory-based approach for lifelong learning in robotic manipulation. Noteworthy, “task-free” does not mean the robot ignores the task context - it should be aware of the instructions combined with the observations to fulfill the job; rather, it means the proposed algorithm effectively handles multiple continually encountered tasks and integrates new knowledge and skills without relying on known task boundaries or explicit task IDs. The concept of “task-free” is properly defined in Section 3. Our method employs a compact storage memory $\mathcal{M}$ that holds a small set of previous tasks’ demonstrations. During training, we replay samples from $\mathcal{M}$ to preserve previously acquired knowledge and skills. Nonetheless, partial forgetting remains inevitable due to the multitasking nature of lifelong learning and limited access to past data. Inspired by human learning mechanisms — where we briefly revisit specific forgotten information to rapidly restore proficiency (Sara, 2000) — we then enable the robot to perform fast local adaptation just before policy deployment. Crucially, we use the same storage memory $\mathcal{M}$ for this adaptation, avoiding any additional storage burden.
Our system retrieves relevant demonstrations based on observation similarity (Du et al., 2023; van Dijk et al., 2024; de Masson D’Autume et al., 2019) and selectively focuses on the most challenging segments where performance degrades—typically the final steps leading to task failure. This automatic selective weighting without relying on human intervention (Spencer et al., 2022; Mandlekar et al., 2020) enables the robots to efficiently restore lost knowledge and skills, facilitating stable, task-free lifelong robot learning. Our key contributions are summarized as:
  • Task-free Retrieval-Based Local Adaptation: A novel local adaptation strategy that retrieves relevant past demonstrations from $\mathcal{M}$ to restore forgotten knowledge and skills, without requiring task IDs or boundaries.
@@
  • Selective Weighting Mechanism: An automatic weighting scheme emphasizing the most challenging segments within the retrieved demonstrations, optimizing real-time adaptation.
  • A General Paradigm for Memory-Based Lifelong Robot Learning: Our approach serves as a plug-and-play solution, complementing existing memory-based lifelong learning algorithms and enabling skill restoration in sequences of open-ended robotics tasks.
2 Related Works
@@ 2.1 Lifelong Robot Learning
Robots operating in continually changing environments need the ability to learn and adapt on-the-fly (Thrun, 1995). In recent years, lifelong robot learning has been applied to SLAM (Yin et al., 2023; Gao et al., 2022; Vödisch et al., 2022), navigation (Kim et al., 2024), and task and motion planning scenarios (Mendez-Mendez et al., 2023). In the context of lifelong learning from demonstrations, robots can 1) acquire skills from non-technical users (Grollman & Jenkins, 2007), or 2) adapt to user-specific task preferences — Chen et al. (2023) proposed a strategy mixture approach to efficiently model new incoming demonstrations, enhancing adaptability.
Furthermore, methods have been developed to improve manipulation capabilities over a robot’s lifespan. Some approaches leverage previous data to facilitate forward transfer but suffer from catastrophic forgetting (Xie & Finn, 2022). Others maintain an expandable skill set to accommodate an increasing number of manipulation tasks (Parakh et al., 2024), or continually update models of manipulable objects for effective reuse (Lu et al., 2022). Large language models have been utilized to improve knowledge transfer (Bärmann et al., 2023; Tziafas & Kasaei, 2024; Wang et al., 2023), and hypernetworks with neural ODEs have been employed to remember long trajectories (Auddy et al., 2023) incrementally. Additionally, (Yang et al., 2022) evaluates how typical supervised lifelong learning methods can be applied in reinforcement learning scenarios for robotic tasks.
To standardize the investigation of lifelong decision-making and bridge research gaps, Liu et al. (2024) introduced LIBERO, a benchmarking platform for lifelong robot manipulation where robots learn multiple atomic manipulation tasks sequentially. Recent works exploring lifelong robot learning based on it include (Liu et al., 2023), which assigns a specific task identity to each task; (Wan et al., 2024), which requires a pre-training phase to build an initial skill set before lifelong learning; and (Lee et al., 2024), tackles multi-stage tasks by incrementally learning skill prototypes for each subgoal, which introduces additional complexities in managing subgoal sequences. However, catastrophic forgetting for lifelong robot learning remains an open challenge, especially when task IDs and boundaries are not available.
2.2 Task-free Lifelong Learning
Despite the success of lifelong learning under clearly labeled task sequences, a significant gap remains in algorithms that can operate independently of task boundaries or IDs during both training and inference, thus aligning more closely with realistic and scalable scenarios. Many approaches (Lee et al., 2020; Chen et al., 2020; Ardywibowo et al., 2022) focus on specialized parameters via expanding network architectures. Meanwhile, researchers have tackled implicit task boundaries in regularization-based methods (Aljundi et al., 2019a; Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017) by consolidating knowledge upon detecting a loss “plateau”. Additionally, (Lässig et al., 2023) introduces a bio-inspired activity-regularization approach employing selective sparsity and recurrent lateral connections, effectively enabling task-free lifelong learning without explicit task boundaries.
Memory-based algorithms further mitigate forgetting by prioritizing informative samples (Sun et al., 2022), discarding less critical examples (Koh et al., 2021), refining decision boundaries (Shim et al., 2021), or enhancing gradient diversity (Aljundi et al., 2019b). Methods aiming to exploit replay buffer in online scenarios (Mai et al., 2021; Caccia et al., 2021) have also demonstrated notable success. However, these algorithms remain largely unexplored in robotic applications that entail sequential decision-making and real-world physical interactions.
@@ 2.3 Robot Learning with Adaptation
Recent advances have shown robots adapting to dynamic environments, such as executing agile flight in strong winds (O’Connell et al., 2022), adapting quadruped locomotion through test-time search (Peng et al., 2020), and generalizing manipulation skills from limited data (Julian et al., 2020). To enable few-shot or one-shot adaptation, meta-learning has been extensively explored (Finn et al., 2017a) and successfully applied to robotics (Kaushik et al., 2020; Nagabandi et al., 2018; Finn et al., 2017b). However, meta-learning methods typically assume access to a full distribution of tasks during meta-training, with both training and testing performed on tasks sampled from this distribution. In contrast, our lifelong robot learning scenario operating sequentially lacks such access, presenting unique challenges of catastrophic forgetting.
@@
Figure 2: Policy Backbone Architecture used in Training and Testing. We input various data modalities into the system, including demonstration images, language descriptions, and the robot arm’s proprioceptive input (joint and gripper states). Pretrained R3M (Nair et al., 2022) and SentenceSimilarity (2024) models process the image and language data respectively. Along with the proprioceptive states processed by an MLP, the embeddings are concatenated and passed through a Transformer to generate temporal embeddings. A GMM (Gaussian Mixture Model) is then used as the policy head to sample actions for the robot. Throughout both training and testing, we utilize a storage memory to store a subset of demonstrations gathered throughout the training process.
3 Preliminaries
To model realistic settings for lifelong robot learning, we define a set of manipulation tasks as $\mathbb{T}=\{\mathcal{T}_{k}\},k=1,2,\dots,T$, where each $\mathcal{T}_{k}$ encompasses a distribution over environmental variations $E_{k}$ (e.g., object positions, robot initial states) and language descriptions $G_{k}$ to guide robot’s actions (e.g., “pick the bottle and put it into the basket,” “place the bottle in the basket please”). From each $\mathcal{T}_{k}$, we sample specific environmental settings $e\sim E_{k}$ and language descriptions $g\sim G_{k}$ to generate a concrete scenario $\mathcal{S}_{n}^{k}\sim p(\mathcal{T}_{k})$, which also serves as the basis for collecting demonstrations $\tau_{n}^{k}$. Multiple demonstrations form the training dataset $\mathcal{D}_{k}=\{\tau^{k}_{n}\},n=1,2,\dots,N$ for task $\mathcal{T}_{k}$.
Notably, multiple tasks may share overlapping distributions in either environmental settings or language descriptions. This natural setting closely mirrors real-world conditions, where it is difficult to determine which task generated a given scenario - tasks are not always divisible. This ambiguity underpins the proposed method’s task-free design, which emphasizes retrieving relevant information rather than relying on task boundaries or IDs.
@@ 3 Preliminaries
The robot utilizes a visuomotor policy learned through behavior cloning to execute manipulation tasks by mapping sensory inputs and language description to motor actions. The policy is trained by minimizing the discrepancy between the predicted actions and expert actions from demonstrations. Specifically, we optimize the following loss function across a sequence of tasks $\mathbb{T}$ with $\mathcal{D}_{k}$. Notably, $\mathcal{D}_{k}$ is only partially accessible for $k<K$ from the storage memory $\mathcal{M}$, where $K$ denotes the current task:
$\theta^{*}$ $=\arg\min_{\theta}\;\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(o_{t},a_{t})\sim\mathcal{D}_{k},\;g\sim G_{k}}$ (1)
$\quad\Biggl{[}\sum_{t=0}^{l_{k}}\mathcal{L}\bigl{(}\pi_{\theta}(o_{\leq t},g),\,a_{t}\bigr{)}\Biggr{]}$
@@
$\theta$ denotes the model parameters, $l_{k}$ represents the number of samples for task $k$, $o_{\leq t}$ denotes the sequence of observations up to time $t$ in demonstration $n$ (i.e., $o_{\leq t}=(o_{0},o_{1},\ldots,o_{t})$), and $a_{t}$ is the expert action at time $t$. The policy output, $\pi_{\theta}(o_{\leq t},g)$, is conditioned on both the observation sequence and the language description.
@@
By optimizing this objective function, the policy effectively continues learning new knowledge and skills in its life span, without the need for task boundaries or IDs, thereby facilitating robust and adaptable task-free lifelong learning.
Figure 3: Data Retrieval. Episodic Memory $\mathcal{M}$ randomly stores a few demonstrations collected during lifelong learning. To retrieve a small number of demonstrations most similar to the current scenario, we compute a weighted distance (Eq (4.1)) using both image and language embeddings. In the storage memory $\mathcal{M}$, red and green circles denote relevant and irrelevant demonstrations, respectively, which include language descriptions, visual observations, expert actions, and joint and gripper states. The retrieved demonstrations are then used for Weighted Local Adaptation.
@@
4 Retrieval-based Weighted Local Adaptation (RWLA) for Lifelong Robot Learning
@@
In this section, we outline our proposed method - Retrieval-based Weighted Local Adaptation (RWLA) - depicted in Figure 1, with corresponding pseudocode in Algorithm 1. To effectively interact with complex physical environments, the network integrates multiple input modalities, including visual inputs from workspace and wrist cameras, proprioceptive inputs of joint and gripper states, and language descriptions.
@@
Instead of training all modules jointly in an end-to-end manner, we employ pretrained visual and language encoders that leverage prior semantic knowledge. Pretrained encoders enhance performance on downstream manipulation tasks (Liu et al., 2023) and are well-suited to differentiate between various scenarios and tasks. They produce consistent representations that are essential both for managing multiple tasks along the lifelong training and for retrieving relevant data to support our proposed local adaptation before policy deployment.
@@
When learning new tasks, the robot preserves previously acquired skills by replaying prior manipulation demonstrations stored in storage memory $\mathcal{M}$ (Chaudhry et al., 2019). Trained with the combined data from the latest scenarios and $\mathcal{M}$, the model can acquire new skills while mitigating catastrophic forgetting of old tasks, thereby maintaining a balance between stability and plasticity (Wang et al., 2024). Figure 2 illustrates the network architecture, and implementation details are provided in Appendix A.2.